OpenStudy (anonymous):

Give an example written in slope intercept form (y = mx + b) perpendicular to the equation y = 2x – 9.. Can you help me with this please?

OpenStudy (anonymous):

the only thing will change is the slope; instead of having y = mx + b, you will have y = -1/m+b in your case m = 2 y = -1/2x -9